• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z. Chai

    Log entry time 23:58:13 on June11,2000

    Entry number 43516

    keyword=Swing Shift Summary

    16:00---Swing shift starts. Finished 3.8C, still need 3.7C more for kin (155,180)
    MCC is fixing the IOC problem in hallA.
    17:24---MCC is sending pulsed beam to Hall A for test.
    17:45---Beam back to Hall A (48uA)
    18:10---EQ3 quenched. Paged Stevens. He is coming to the Hall to check it.
    Reboot the target IOC.
    19:18---Stevens finished checking and resetting the EQ3. Starts to cycle the
    EQ3.
    19:36---MCC is ready to send the beam to Hall A.
    19:40---Beam is back (47uA)
    20:16---EQ3 quenched again. Paged Stevens. He told us to reset it. Adam went down to the Hall to reset the EQ3.
    20:54---Beam is back(46.4uA)
    21:18---EQ3 quenched. Mark Stevens is coming to check it.
    22:45---Beam back to Hall A
    23:43---BCM tripped.
    23:49---Beam back.

    Time........Run#......Evts.........Current....Charge....Comment
    17:45........1952........1.3M.......47.5uA.....61mC......Good
    18:10........1953........0.5M.......47.2uA.....24mC......EQ3 quench
    19:40........1954........1.3M.......46.8uA.....62mC......Good
    20:03........1955........0.7M.......46.8uA.....31mC......EQ3 quench
    20:54........1956........1.3M.......46.5uA.....59mC......EQ3 quench
    21:18........1957.........................................................Short junk run
    22:47........1958........1.3M.......48.7uA......60mC......Good
    23:10........1959........1.3M.......48.7uA......60mC......Good
    23:33........1960........0.5M.......48uA.........23mC......BCM tripped.
    23:49........1961........